Jancie has $180,000 to invest to obtain annual income. She wants some of it invested in safe Certificates of Deposit yielding 5% The rest she wants to invest in AA bonds yielding 12% per year. How much should she invest in each to realize exactly $17,400 per year?


Answer by Boreal(15235)

x at 5%
180000-x at 12%
.05x+.12(180000-x)=17400
.05x+21600-.12x=17400
-.07x=4200
x=$60000@5% for $3000
180000-x=$120000@12%=$14400
They add to $17,400